1. A privately owned summer camp for youngsters hasthe followingdata for a 12-weeksession:
Charge per camper $120 per week
Fixed costs $48,000 per session
Variable cost per camper $80 per week
Capacity 200 campers

(a) Develop the mathematical relationships for totalcost and total revenue.
(b) What is the total number of campers that willallow the camp to just break even?
(c) What is the profit or loss for the 12-week sessionif the camp operates at 80% capacity?
1
Expert's answer
2020-05-08T14:18:07-0400

(a) "total cost=Fixed cost+Variable cost =48000+80\times number of campers\times12"

"total revenue=Capacity campers *Charge per camper\times number of campers=120\times 12\times number of campers"

(b) TR=TC

"48000+80\times12\times number= 120\times12\times number"

"48000= 480\times number"

number=100

(c)"200\times0.8=160" number of campers

"120\times160\times12-48 000-160\times80\times12=230 400-48 000-153 600=28 800" profit